OpenStudy (anonymous):

solve by elimination 0.3x-0.2y=4 0.4x+0.3y=-1

OpenStudy (amistre64):

muliply by 10 to get rid of the decimals... that will make it look less intimidating

OpenStudy (amistre64):

3x-2y = 40 4x +3y = -10

OpenStudy (yuki):

if you multiply 40 on the first eqn and 30 on the second, you can see which ones can be eliminated

OpenStudy (yuki):

your idea is to make the coefficients of either x or y to be equal

OpenStudy (yuki):

for most elimination problems, that will work

OpenStudy (amistre64):

to eliminate the x variables; *4 the top and *3 the bottom...then subtract

OpenStudy (amistre64):

12x -8y = 160 12x +9y = -30

OpenStudy (amistre64):

we could combine these steps by mulitply one equation by a negative instead... 12x -8y = 160 -12x -9y = 30 --------------- -17y = 190 y = -190/17

OpenStudy (anonymous):

you lost me

OpenStudy (amistre64):

where?

OpenStudy (anonymous):

12x-8y=160 12x-9y=-30 how did you get those

OpenStudy (amistre64):

by subtraction...... 12x -8y = 160 12x -8y = 160 -(12x +9y = -30) means -12x -9y = 30 ---------------- ----------------

OpenStudy (amistre64):

tell me: does this equal out? 5 5 -(-4) +4 ---- ---- ???

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

thats all i did..... just with more numbers involved

OpenStudy (anonymous):

how did you get the 12x -8y = 160 & -(12x +9y = -30)

OpenStudy (amistre64):

elimination means 'get rid of something'...right? the only way to get rid of something; to eliminate it; is to get it to zero... otherwise there is something still there right?

OpenStudy (anonymous):

right

OpenStudy (amistre64):

you are given 2 equations and are asked to eliminate something.... that would either be the xs or the ys right? I choose the xs cause they were in front.

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

do you recall that in an equation..whatever you do to one side has to be done to the other to keep it equal?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

5+3 = 8 when we multiply this by 4 is it still equal? 4(5+3 = 8)..... 20 + 12 = 32 ???

OpenStudy (anonymous):

right

OpenStudy (amistre64):

do the variables make a difference to it? 4(5x + 3y = 4)......20x +12y = 16...still equal??

OpenStudy (amistre64):

we are not changing the 'value'...just the form, its shape, the way it looks..... the value remains the same. right?

OpenStudy (anonymous):

right

OpenStudy (amistre64):

so lets use this to our advantage.. we want some part of your equation to cancel each other out... what wer ethe equations again?

OpenStudy (amistre64):

0.3x-0.2y=4 0.4x+0.3y=-1 we can multiply both equation by 10 to get rid of the decimal and make it easier to look at right? we change the shape; but not the value

OpenStudy (anonymous):

k

OpenStudy (amistre64):

3x - 2y = 40 4x +3y = -10 we good here?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

we have no terms that are the same...so we have to modify some more..... to get rid of one of them.....or rather a set of them.

OpenStudy (amistre64):

if we pick 'x'...what value do 4x and 3x have in common?

OpenStudy (anonymous):

x?

OpenStudy (amistre64):

x is not a 'value' :) 4x and 3x..... do they have 12x in common? 8x? 29x?

OpenStudy (amistre64):

x is a valued pick; but that makes things messier lol

OpenStudy (amistre64):

valid pick is what its spose to say...

OpenStudy (amistre64):

4x(3) = 12x 3x(4) = 12x.....right?

OpenStudy (anonymous):

right

OpenStudy (amistre64):

so the top can be *4; and the bottom *3 4(3x - 2y = 40) 12x -8x =160 3(4x +3y = -10) 12x +9y = -30 havent changed anything but the way they look correct?

OpenStudy (anonymous):

correct

OpenStudy (amistre64):

but if we add these together...nothing is eliminated is it?... what do we have to do to get an elimination now?

OpenStudy (anonymous):

wouldnt x be eliminated

OpenStudy (amistre64):

how? 12x + 12x = 24x..... that aint zero..

OpenStudy (anonymous):

your right i was thinking bout -12x +12x continue plz

OpenStudy (amistre64):

thats right... we need to change the way one of them looks again to get a -12x... 12x -8x =160 12x -8x =160 -1(12x +9y = -30) -12x -9y = 30..now were good to go

OpenStudy (amistre64):

can we eliminate now?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

we get: -17y = 190..... solve for y y = -190/17. right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

since we have a solid value for y...use it to find x

OpenStudy (anonymous):

wait how does 190 become -190 and -17 becomes 17

OpenStudy (amistre64):

the negative doesnt care what its attached to..... any (+)(-) combonation results in a (-)...right? does 190/-17 equal something different than -190/17?

OpenStudy (anonymous):

i guess not

OpenStudy (amistre64):

the proper ettiquiette is to put the sign to the top...

OpenStudy (amistre64):

but in some cases its good to know that you can move it around

OpenStudy (amistre64):

what wouldve hapened aif we combined steps?

OpenStudy (amistre64):

.4(.3x -.2y = 4) .12x -.08y = 1.6 -.3(.4x+.3y =-1) -.12x -.09y = .3 ----------------- -.17y = 1.9 y = -1.9/.17 = -190/17

OpenStudy (anonymous):

right

OpenStudy (amistre64):

0.4x+0.3y=-1 .4x + .3(-190/17) = -1....solve for x

OpenStudy (anonymous):

-5.7

OpenStudy (anonymous):

or is it -133/17

OpenStudy (amistre64):

4x = -10 + 570/17 4x = (-170 + 570)/17 4x = 400/17 x = 400/4.17 = 100/17

OpenStudy (amistre64):

4.17 is a lazy way of writing 4*17

OpenStudy (anonymous):

how did yiu get 570/17

OpenStudy (amistre64):

i used 3(-190/17) instead of .3(-190/17) 3(-190) = -570 ------ ------ 17 17 4x - 570/17 = -10 4x = -10 + (570/17)...

OpenStudy (anonymous):

oh ok i was doing it by .3

OpenStudy (amistre64):

:) .3 was fine, but you would have had to change the decimal eventually

OpenStudy (amistre64):

1/.17 cahnges to 100/17

OpenStudy (anonymous):

-10*17=-170

OpenStudy (amistre64):

.4x + .3(-190/17) = -1 .4x - 57/17 = -1 .4x = -1 + (57/17) .4x = (-17+57)/17 .4x = 40/17 x = 40/(.4*17) x = 400/(4*17) x = 100/17

OpenStudy (anonymous):

ok i think i got it

OpenStudy (amistre64):

i think i see you trying to get rid of the 17 on the bottom :)

OpenStudy (anonymous):

i just dont know how to do this stuff

OpenStudy (anonymous):

taking midterms this week these questions are for the practice exam

OpenStudy (amistre64):

68x - 570 = -170 68x = -170 + 570 x = 400/68 x = 400/4 100 ------- = ----- 68/4 17

OpenStudy (amistre64):

i got a final grade of 89.06%...and it gave me an 'A' for the course ;)

OpenStudy (amistre64):

i dropped 3 writing assignments and still passed with an 'a' lol

OpenStudy (anonymous):

thats good! right now i have a 77.37%

OpenStudy (anonymous):

how did you get those answers above 68x....

OpenStudy (amistre64):

i times everything by 17 to get rid of the fraction...

OpenStudy (amistre64):

4(17) = 68

OpenStudy (anonymous):

right but 3(17)=51

OpenStudy (amistre64):

3(-190)(17) ---------- = 3(-190) (17)

OpenStudy (amistre64):

17/17 = 1

OpenStudy (anonymous):

x=400/68 i divide that by?

OpenStudy (amistre64):

thats why its called getting rid of a fraction ;)

OpenStudy (amistre64):

400/68 has 4 in common..remember 4(17)? and 4(100) = 400

OpenStudy (amistre64):

divide them both by 4

OpenStudy (anonymous):

ok now that way makes sense to me

OpenStudy (amistre64):

you seem to see numbers and math is a solid concrete building that cant be moved or messes with. but its really more like silly putty...bend it to your will :)

OpenStudy (amistre64):

algebra has 5 rules: and only 5 rules...

OpenStudy (anonymous):

whats that?

OpenStudy (amistre64):

1. numbers can be moved around (commutative) 2. numbers canbe grouped in different ways (associative) 3. every number has an evil twin (inverse). 4. there is one number that changes nothing (identity) 5. distribute numbers thru paranthesis (distributive)

OpenStudy (anonymous):

i will have to make a note of those

OpenStudy (amistre64):

3 and 4 are split in 2 by some people to get an additivie inverse/multiplicative inverse and an additive indentity/multiplicative identity

OpenStudy (amistre64):

tell me; the 1st rule says we can move number around... 4-5 is the same as -5+4 right?